Assume the $19,500 Treasury bill, 4% for 15 weeks. Calculate the effective rate of interest. (Use calendar year. Round your answer to the nearest hundredth percent.) Effective rate of interest:
Answer #1

EAR = [1 + r]m - 1

= [1.04](52/15) - 1 = 1.1456 - 1 = 0.1456, or 14.56%

m = No. of compounding periods in a year